Peter Pine is doing his budget. He discovers that he has spent an average of $225.00 a month on entertainment for the year with a standard deviation of $75.00. What percent of his monthly expenses for this category would he expect to fall between $60.00 and $390.00?
The z for $60.00 = - a0.
The percent of area associated with $60.00 = a1 %.
The z for $390.00 = a2.
The percent of area associated with $390.00 = a3 %.
Adding the two percentages together, Peter calculates his answer to be a4 %.

Answer by stanbon(75887)

Peter Pine is doing his budget. He discovers that he has spent an average of $225.00 a month on entertainment for the year with a standard deviation of $75.00. What percent of his monthly expenses for this category would he expect to fall between $60.00 and $390.00?
The z for $60.00 = (60-225)/75 = -2.2
------------------------------------------
The percent of area between z = 0 and z = -2.2 = 0.4861
-----
The z for $390.00 = 2.2
The percent of area associated with $390.00 = +0.4861
--------------------
Adding the two percentages together,
Peter calculates his answer to be 2*0.4861 = 0.9722
=================
Cheers,
Stan H.
=======================